# Safe Output Regulation of Coupled Hyperbolic PDE-ODE Systems

### v1

@@ Abstract
This paper presents a safe output regulation control strategy for a class of systems modeled by a coupled $2\times 2$ hyperbolic PDE-ODE structure, subject to fully distributed disturbances throughout the system. A state-feedback controller is developed by the nonovershooting backstepping method to simultaneously achieve exponential output regulation and enforce safety constraints on the system output, which is the state furthest from the control input. To handle unmeasured PDE states and external disturbances, a state observer and a disturbance estimator are designed. Explicit bounds on the estimation errors are derived and used to construct a robust safe regulator that accounts for the uncertainties. The proposed control scheme guarantees that: 1) If the system output is initially within the safe region, it remains there; otherwise, it will be rescued to the safety within a prescribed time; 2) The output tracking error converges to zero exponentially; 3) The observer accurately estimates both the distributed states and external disturbances, with estimation errors converging to zero exponentially; 4) All signals in the closed-loop system remain bounded. The effectiveness of the proposed method is demonstrated through a UAV delivery scenario with a cable-suspended payload, where the payload is regulated to track a desired reference while avoiding collisions with barriers.
@@ 1.1 Output Regulation of Coupled Hyperbolic PDEs
The output regulation problem seeks to design the control law that ensures the system output tracks desired references and/or rejects undesired disturbances. In recent years, considerable attention has been devoted to output regulation for infinite-dimensional systems governed by partial differential equations (PDEs). Among various methodologies, the backstepping approach has emerged as a widely adopted and powerful framework, whose foundational development was established in [14]. Upon the backstepping-based boundary stabilization of $2\times 2$ hyperbolic PDEs presented in [20] and [8], initial contributions toward backstepping-based output regulation for $2\times 2$ hyperbolic PDEs appeared in [1]. Subsequent extensions to more general heterodirectional transport PDEs were proposed in [2], [11], and [27]. For wave-type PDEs, similar regulator designs were developed in [28] and [29]. Research efforts were also directed at PDE-ODE interconnected systems, which naturally arise in many applications. Backstepping boundary control designs for interconnected $2\times 2$ hyperbolic PDEs and ODEs were proposed in [17], and the ODE-PDE-ODE configurations were further addressed in [23], [21], [9], and [7]. Subsequently, the output regulation of hyperbolic PDEs sandwiched between two ODEs was explored in [31], and the regulation problems involving coupled wave PDE-ODE systems and hyperbolic PDEs coupled with nonlinear ODEs were investigated in [12] and [30], respectively.
Despite these advances, most existing works only focus on stabilization and output regulation goals without addressing safety constraints—that is, ensuring that the system output state remains within a predefined safe region during the regulating operation. In many applications, such as autonomous vehicles, robotics, and UAV systems, where collision avoidance is essential, safety is critical. Therefore, integrating output regulation objectives with safety guarantees for PDE systems, i.e., safe output regulation of PDEs, remains an open and significant research topic.
@@ 1.2 Safe Control
However, safe control of PDE systems remains largely underexplored. The first result in this area was presented in [13], which proposed a CBF-based boundary control scheme for a Stefan model involving parabolic PDEs with actuator dynamics. The safe regulation of the hyperbolic PDEs was proposed in [24]. Yet, these approaches assume full-state availability and do not address output-feedback scenarios with external disturbances. Designing observer-based output-feedback safety controllers is challenging even for disturbed ODE systems, as ensuring safety in the presence of state estimation errors and external disturbances requires significant enhancements to the robustness of the CBF framework. These challenges become even more pronounced in PDE systems, making the development of output-feedback safe PDE control a highly nontrivial and largely open research problem.
@@ 1.4 Notation
  • We use the notation $L^{2}(0,1)$ for the standard space of the equivalence class of square-integrable, measurable functions $f:(0,1)\to\mathbb{R}$, with $\|f\|^{2}:=\int_{0}^{1}f(x)^{2}dx<+\infty$ for $f\in L^{2}(0,1)$.
@@ 2.1 Plant
$\forall(x,t)\in[0,1]\times[0,\infty)$, where the function $U(t)$ is the control input to be designed. The scalars $z(x,t)\in\mathbb{R},w(x,t)\in\mathbb{R}$ are states of the PDEs, $Y^{T}(t)=[y_{1},y_{2},\cdots,y_{n}]\in\mathbb{R}^{n}$ are ODE states. The disturbances $d(t)\in\mathbb{R}^{m_{d}\times 1}$ are unmeasurable. The plant parameters $d_{1},d_{2}$, the transport speed $q_{1}>0,q_{2}>0$, and the matrices $G_{1}\in\mathbb{R}^{n\times m_{d}},G_{2},G_{3}\in(C[0,1])^{1\times m_{d}}$ regarding the disturbance input locations, are known and arbitrary. The subsystem $Y$-ODE is strict-feedback linear, where the matrix $A$, the column vector $B$ are in the form of
@@
and $B=(0,0,\cdots,b)^{T}$, with arbitrary constants $a_{i,j}$, and $b>0$ (without any loss of generality for $b<0$). This indicates that the $Y$-ODE is in the controllable form which covers many practical models. The matrix $C\in\mathbb{R}^{1\times n}$ is arbitrary satisfying that $(A,C)$ is observable.
@@ 2.2 Exogenous signal model
where the spectrum $\sigma(S)$ of the known diagonalizable matrix $S\in\mathbb{R}^{n_{r}\times n_{r}}$ only contains eigenvalues on the imaginary axis, which allows the modelling of bounded and persistently acting exogenous signals. In particular, this signal model can generate the exogenous signals as constant or trigonometric functions of time as well as linear combinations of both signal forms. Defining $S={\rm bdiag}(S_{r},S_{d})$ with $S_{r}\in\mathbb{R}^{n_{r}\times n_{r}}$, $S_{d}\in\mathbb{R}^{n_{d}\times n_{d}}$, and $v={\rm col}(v_{r},v_{d})$ where $v_{r}\in\mathbb{R}^{n_{r}}$, $v_{d}\in\mathbb{R}^{n_{d}}$, one obtains the reference model and the disturbance model, respectively,
@@ 2.3 Control Objective
The output tracking error $e(t)=y_{1}(t)-r(t)$ of the system converges to zero while staying in the safe region in the sense that the time-varying barrier function $h(e(t),t)$ is non-negative, i.e., $h(e(t),t)\geq 0$. Because of the delay property of the transport PDE, there is no control for the distal ODE in the time interval $t_{0},\bar{t}_{0}]$, where $t_{0}$ is the initial time and $\bar{t}_{0}$ is the time when the control action begins to regulate the ODE, i.e.,
$\bar{t}_{0}=t_{0}+\frac{1}{q_{2}}$ (16)
namely, the initial regulation time. We impose the following assumption regarding the time-varying barrier function $h$ for the output signal.
@@
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\{\ell\in\mathbb{R}|h(\ell,t)\geq 0\}$, $t\in[\bar{t}_{0},\infty)$ when $h$ is positive at $t=\bar{t}_{0}$, or otherwise $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[\bar{t}_{0},\infty)$. Besides, $|h(e(t),t)|<\infty\Rightarrow|e(t)|<\infty$ and $\lim_{t\to\infty}h(e(t),t)=0\Rightarrow\lim_{t\to\infty}e(t)=0.$
In Assumption 2, the fact $\lim_{t\to\infty}h(e(t),t)=0\Rightarrow\lim_{t\to\infty}e(t)=0$ implies the prescribed safe region converges to the desired reference as $t\to\infty$, which is reasonable in the output regulation problem.
@@ 3.1 First transformation
and the system matrix $A_{\rm z}$ represents a chain-of-integrators which will be helpful for advancing to high-relative-order barrier functions in the next transformation. The details of the first transformation are shown in Appendix 0.1.
@@ 3.2 Second transformation
where $t_{a}$ and $\epsilon$ are arbitrarily positive design parameters. The function $\sigma(t)$ is designed to address scenarios where the states are in the unsafe region at the initial regulation time $\bar{t}_{0}$. Specifically, when the initial state $e(\bar{t}_{0})$ fall outside the original safe region—indicated by the condition $(h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$—a new barrier function is created to guide the state back to the original safe region. The constant $\epsilon$ indicates the distance from the safe barrier of the initial value of the new barrier function $h_{1}$, and $t_{a}$ is the upper bound of the time taken to rescue to the safe region. We only ensure the state, which falls into the unsafe region at the initial regulation time $\bar{t}_{0}$, to rescue to safety within a prescribed time $t_{a}$, but not prescribed-time regulation to the safe boundary. The evolution of $e(t)$ on $t\in[t_{0},\bar{t}_{0}]$ can be expressed as the initial states as
@@
$\forall a\in[0,\frac{1}{q_{2}}]$, where $Z(t_{0}+a)$ is defined by (B.4) in Appendix 0.2, which is determined by the initial states $Y(t_{0}),w[t_{0}],z[t_{0}]$ and $v(t_{0})$. Therefore, $h(e(\bar{t}_{0}),\bar{t}_{0})$ can be determined by the initial states of the plant according to (50), denoted as
@@
because of $z_{1}(t)=e(t)$ (46). Considering the fact $\frac{\partial h_{n}}{\partial z_{n}}=\frac{\partial h_{n-1}}{\partial z_{n-1}}=\frac{\partial h_{n-2}}{\partial z_{n-2}}=\cdots=\frac{\partial h_{1}}{\partial z_{1}}=\frac{\partial h}{\partial e}=\vartheta$ that is obtained from (48), using (48) for $i=n$, we have $\dot{h}_{n}(\underline{z}_{n},t)+k_{n}h_{n}=b(\vartheta w(0,t)+f(\underline{y}_{n}(t),v(t),t)-\vartheta\frac{1}{b}(\lambda_{n}+P_{r}S^{n})v(t)+\vartheta K^{T}Y(t))$, where
@@ 3.3 Third transformation
In order to remove the in-domain coupling destabilizing terms from the $2\times 2$ hyperbolic PDE system and cancel the extra terms in (37), we propose the following backstepping transformations:
@@
$-\int_{0}^{x}{\varphi}(x,y)w(y,t)dy-\gamma(x)Y(t)-\bar{\gamma}(x)v(t)-q(x,t),$ (62)
@@
$-\int_{0}^{x}{\Phi}(x,y)w(y,t)dy-\lambda(x)Y(t)-\bar{\lambda}(x)v(t)-p(x,t)$ (63)
where ${\phi},\varphi,\gamma,\bar{\gamma},\Psi,\Phi,\lambda,\bar{\lambda}$ are defined in Appendix 0.3, and where $p(x,t)$ and $q(x,t)$ satisfy
$p_{t}(x,t)=q_{2}p_{x}(x,t),~~p(0,t)=\frac{-1}{\vartheta(z_{1}(t),t)}f(\underline{z}_{n},t),$ (64)
$q_{t}(x,t)=-q_{1}q_{x}(x,t),~~q(0,t)=p*p(0,t).$ (65)
The solution $p$ is
$p(x,t)=\frac{-1}{\vartheta(z_{1}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}})}f\left(\underline{z}_{n}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}}\right)$
@@
$\alpha(0,t)=$ $p\beta(0,t),$ (68)
${\alpha_{t}}(x,t)=$ $-{q_{1}}{\alpha_{x}}(x,t),$ (69)
${\beta_{t}}(x,t)=$ ${q_{2}}{\beta_{x}}(x,t),$ (70)
$\beta(1,t)=$ $0$ (71)
@@
$-(\acute{G}_{5}-\bar{\lambda}(1)){v}(t)+p(1,t)$ (72)
where $p(1,t)$ is given by (47)–(49), (53), (3.3), (B.4), ${\Psi}(1,y),{\Phi}(1,y)$, $\lambda(1)$ are determined by (C.16)–(C.21), and $\bar{\lambda}(1)$ is given by (C.12), (C.13), that is,
$p(1,t)=$ $\frac{-f(\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}{\vartheta(C_{1}\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}$

### v2

@@ Abstract
This paper presents a safe output regulation control strategy for a class of systems modeled by a coupled $2\times 2$ hyperbolic PDE-ODE structure, subject to fully distributed disturbances throughout the system. A state-feedback controller is developed by the nonovershooting backstepping method to simultaneously achieve exponential output regulation and enforce safety constraints on the regulated output that is the state furthest from the control input. To handle unmeasurable states and external disturbances, a state observer and a disturbance estimator are designed. Explicit bounds on the estimation errors are derived and used to construct a robust safe regulator that accounts for the uncertainties. The proposed control scheme guarantees that: 1) If the regulated output is initially within the safe region, it remains there; otherwise, it will be rescued to the safety within a prescribed time; 2) The output tracking error converges to zero exponentially; 3) The observer accurately estimates both the distributed states and external disturbances, with estimation errors converging to zero exponentially; 4) All signals in the closed-loop system remain bounded. The effectiveness of the proposed method is demonstrated through a UAV delivery scenario with a cable-suspended payload, where the payload is regulated to track a desired reference while avoiding collisions with barriers.
@@ 1.1 Output Regulation of Coupled Hyperbolic PDEs
The output regulation problem seeks to design the control law that ensures the system output tracks desired references and/or rejects undesired disturbances. In recent years, considerable attention has been devoted to output regulation for infinite-dimensional systems governed by partial differential equations (PDEs). The Internal Model Principle (IMP), serving as a fundamental theoretical framework for output regulation, whose primary advantage lies in its inherent robustness against internal uncertainties, has been extensively investigated for infinite-dimensional systems in [18, 19, 20, 21].
In addition, the integration of output regulation theory with the backstepping methodology [14] has led to systematic and constructive procedures for regulator design in PDE systems. Upon the backstepping-based boundary stabilization of $2\times 2$ hyperbolic PDEs presented in [24] and [8], initial contributions toward backstepping-based output regulation for $2\times 2$ hyperbolic PDEs appeared in [1]. Subsequent extensions to more general heterodirectional transport PDEs were proposed in [2], [11], and [33]. For wave-type PDEs, similar regulator designs were developed in [34] and [35]. Research efforts were also directed at PDE-ODE interconnected systems, which naturally arise in many applications. Backstepping boundary control designs for interconnected $2\times 2$ hyperbolic PDEs and ODEs were proposed in [17], and the ODE-PDE-ODE configurations were further addressed in [28], [26], [9], and [7]. Subsequently, the output regulation of hyperbolic PDEs sandwiched between two ODEs was explored in [37], and the regulation problems involving coupled wave PDE-ODE systems and hyperbolic PDEs coupled with nonlinear ODEs were investigated in [12] and [36], respectively. A more detailed literature review of backstepping output regulation/disturbance rejection in PDEs can be found in Sec. 3.5 of [25]. Despite progress in PDE output regulation, guaranteed safety remains unaddressed. Ensuring that outputs remain within safe operating regions is crucial for autonomous systems, thus the problem of safe output regulation of PDEs is essential and open.
@@ 1.2 Safe Control
However, safe control of PDE systems remains largely underexplored. The first result in this area was presented in [13], which proposed a CBF-based boundary control scheme for a Stefan model involving parabolic PDEs with actuator dynamics. Adaptive safe control for hyperbolic PDEs was proposed in [30]. Yet, these approaches address full-state control but do not address output regulation in the presence of unmeasured states and external disturbances.
@@ 1.4 Notation
  • We use the notation $L^{\infty}(0,1)$ for the standard Banach space of essentially bounded, measurable functions $f:(0,1)\to\mathbb{R}$, equipped with the norm $\|f\|_{\infty}:=\text{ess sup}_{x\in(0,1)}|f(x)|<+\infty$ for $f\in L^{\infty}(0,1)$.
@@ 2.1 Plant
$\forall(x,t)\in[0,1]\times[0,\infty)$, where the function $U(t)$ is the control input to be designed. The scalars $z(x,t)\in\mathbb{R},w(x,t)\in\mathbb{R}$ are states of the PDEs, $Y^{T}(t)=[y_{1},y_{2},\cdots,y_{n}]\in\mathbb{R}^{n}$ are ODE states. The disturbances $d(t)\in\mathbb{R}^{m_{d}\times 1}$ are unmeasurable. The plant parameters $d_{1},d_{2}$, $q$, the transport speed $q_{1}>0,q_{2}>0$, the matrices $G_{1}\in\mathbb{R}^{n\times m_{d}},G_{2},G_{3}\in(C[0,1])^{1\times m_{d}}$ regarding the disturbance input locations, and $p\neq 0$, are known and arbitrary. The subsystem $Y$-ODE is strict-feedback linear, where the matrix $A$, the column vector $B$ are in the form of
@@
and $B=(0,0,\cdots,b)^{T}$, with arbitrary constants $a_{i,j}$, and $b>0$ (without any loss of generality for $b<0$). This indicates that the $Y$-ODE is in the controllable form, which covers many practical models. The matrix $C\in\mathbb{R}^{1\times n}$ is arbitrary.
@@ 2.2 Exogenous signal model
where the spectrum $\sigma(S)$ of the known diagonalizable matrix $S\in\mathbb{R}^{n_{v}\times n_{v}}$ only contains eigenvalues on the imaginary axis, which allows the modelling of bounded and persistently acting exogenous signals. In particular, this signal model can generate the exogenous signals as constant or trigonometric functions of time as well as linear combinations of both signal forms. Defining $S={\rm bdiag}(S_{r},S_{d})$ with $S_{r}\in\mathbb{R}^{n_{r}\times n_{r}}$, $S_{d}\in\mathbb{R}^{n_{d}\times n_{d}}$, and $v={\rm col}(v_{r},v_{d})$ where $v_{r}\in\mathbb{R}^{n_{r}}$, $v_{d}\in\mathbb{R}^{n_{d}}$, one obtains the reference model and the disturbance model, respectively,
@@ 2.3 Control Objective
There are two measurable outputs in the considered plant, one is $z(1,t)$ that will be utilized in the extended observer for both unmeasurable states and exogenous signals, and the other one is $y_{1}=C_{1}Y(t)$ that is the output to be safely regulated, which is described below. The vector $C_{1}=[1,0,\cdots,0]$ satisfies the following assumption.
@@
The pair $(A,C_{1})$ is observable.
The control objective is to ensure that the output tracking error $e(t)=y_{1}(t)-r(t)$ of the system converges to zero while staying in the time-varying safe region in the sense that $e(t)\in\mathcal{C}(t),\forall t\geq\bar{t}_{0}$, where $\mathcal{C}(t)$ is a time-dependent safe set $\mathcal{C}(t):=\{e\in\mathbb{R}\mid h(e,t)\geq 0\}$ defined by the time-varying barrier function $h(e(t),t)$. Focusing on $y_{1}$ is only for ease of exposition and does not restrict the applicability of the result. The result is not applicable only to systems whose first state only is constrained. The result is applicable to all systems diffeomorphically transformable into the form (1), (11), which is equivalent to the input-output feedback linearization-like treatment with high-order CBFs (HOCBFs). And the result is also applicable, with an observer, to all systems transformable into (1) with a transformation acting not only on the measured state but on the entire (unmeasured) state. An example is given in the following remark.
Remark 1.
The control design is applicable to the case where the distal ODE subsystems (denoted as the XX–ODE) are described by general matrices $\bar{A}\in\mathbb{R}^{n},\bar{B}\in\mathbb{R}^{n\times 1},\bar{C}\in\mathbb{R}^{1\times n}$, assuming that the pair $(\bar{A},\bar{C})$ is observable and that the XX–ODE has relative degree $n$. Such systems are also covered by the framework considered in this paper. Indeed, the XX–ODE can be transformed into the YY–ODE studied here via the coordinate transformation $Y(t)=RX(t)$, $R=[\bar{C}^{T},(\bar{C}\bar{A})^{T},\ldots,(\bar{C}\bar{A}^{n-1})^{T}]^{T}$. Since $(\bar{A},\bar{C})$ is observable, the matrix $R$ is nonsingular. The relative degree $n$ that implies that $\bar{C}\bar{A}^{i}\bar{B}=0$ for $i=0,\cdots,n-2$ and $\bar{C}\bar{A}^{n-1}\bar{B}\neq 0$. Then it is ensured that the transformed system’s system matrix, which is in the companion (observable canonical) form whose last row consists of the coefficients of the characteristic polynomial of $\bar{A}$ according to Hamilton-Cayley theorem, is covered by the structure $A$ in (11), and the resulting input and output matrices also match $B,C_{1}$ matrices considered in this paper. Furthermore, the system output is preserved under the transformation, namely, $\bar{C}X=C_{1}Y=y_{1}$. Therefore, the safe regulation of the YY–ODE output achieved by the following control design directly implies the safe regulation of the original XX–ODE output.
The safety goal is now to ensure $h(e(t),t)\geq 0$ from the initial regulation time:
$\bar{t}_{0}=t_{0}+\frac{1}{q_{2}}.$ (16)
The introduction of initial regulation time $\bar{t}_{0}$ in (16) is due to the delay property of the transport PDE. Specifically, there is no control influence on the distal ODE over the time interval $t_{0},\bar{t}_{0}]$, where $t_{0}$ is the initial time. We impose the following assumption regarding the time-varying barrier function $h$ for the output signal.
Assumption 3.
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\{\ell\in\mathbb{R}|h(\ell,t)\geq 0\}$, $t\in[\bar{t}_{0},\infty)$ when $h$ is positive at $t=\bar{t}_{0}$, or otherwise $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[\bar{t}_{0},\infty)$. Besides, there exists a positive constant $\xi$ such that $|e(t)|<\xi|h(e(t),t)|$
Assumption 3 implies that $\lim_{t\to\infty}h(e(t),t)=0\Rightarrow\lim_{t\to\infty}e(t)=0$, which means the prescribed safe region converges to the desired reference as $t\to\infty$, which is reasonable in the output regulation problem.
In the derivations below, we work under the standing strong assumption that the plant, and later the corresponding closed-loop and observer systems, admit classical solutions on the time interval under consideration, so that all temporal and spatial derivatives appearing in the PDEs and all boundary traces used in the control law and proofs are well defined pointwise.
@@ 3.1 First transformation
and the system matrix $A_{\rm z}$ represents a chain-of-integrators which will be helpful for advancing to high-relative-order barrier functions in the next transformation. The details of the first transformation are shown in Appendix A.
@@ 3.2 Second transformation
where $t_{a}$ and $\epsilon$ are arbitrarily positive design parameters. The function $\sigma(t)$ is designed to address scenarios where the states are in the unsafe region at the initial regulation time $\bar{t}_{0}$. Specifically, when the initial state $e(\bar{t}_{0})$ fall outside the original safe region—indicated by the condition $(h(e(\bar{t}_{0}),\bar{t}_{0})\leq 0$—a new barrier function is created to guide the state back to the original safe region. The constant $\epsilon$ indicates the distance from the safe barrier of the initial value of the new barrier function $h_{1}$, and $t_{a}$ is the upper bound of the time taken to rescue to the safe region. We only ensure the state, which falls into the unsafe region at the initial regulation time $\bar{t}_{0}$, to rescue to safety within a prescribed time $t_{a}$, but not prescribed-time regulation to the safe boundary. Please note that even though $t_{a}$ can be assigned arbitrarily, a smaller value of $t_{a}$ leads to larger control gains, and thus a proper trade-off should be considered in practical implementation. The evolution of $e(t)$ on $t\in[t_{0},\bar{t}_{0}]$ can be expressed as the initial states as
@@
$\forall a\in[0,\frac{1}{q_{2}}]$, where $Z(t_{0}+a)$ is defined by (B.5) in Appendix B, which is determined by the initial states $Y(t_{0}),w[t_{0}],z[t_{0}]$ and $v(t_{0})$. Therefore, $h(e(\bar{t}_{0}),\bar{t}_{0})$ can be determined by the initial states of the plant according to (50), denoted as
@@
because of $z_{1}(t)=e(t)$ (46). Considering the fact
$\frac{\partial h_{n}}{\partial z_{n}}=\frac{\partial h_{n-1}}{\partial z_{n-1}}=\frac{\partial h_{n-2}}{\partial z_{n-2}}=\cdots=\frac{\partial h_{1}}{\partial z_{1}}=\frac{\partial h}{\partial e}=\vartheta$ (53)
that is obtained from (48), using (48) for $i=n$, we have $\dot{h}_{n}(\underline{z}_{n},t)+k_{n}h_{n}=b(\vartheta w(0,t)+f(\underline{y}_{n}(t),v(t),t)-\vartheta\frac{1}{b}(\lambda_{n}+P_{r}S^{n})v(t)+\vartheta K^{T}Y(t))$, where
@@ 3.3 Third transformation
In order to remove the in-domain coupling destabilizing terms from the $2\times 2$ hyperbolic PDE system, the extra terms in (37), and the dependence on the signal model state $v(t)$, we propose the following backstepping transformations:
@@
$-\int_{0}^{x}{\varphi}(x,y)w(y,t)dy-\gamma(x)Y(t)-\bar{\gamma}(x)v(t)-\bar{\varsigma}(x,t),$ (63)
@@
$-\int_{0}^{x}{\Phi}(x,y)w(y,t)dy-\lambda(x)Y(t)-\bar{\lambda}(x)v(t)-{\varsigma}(x,t)$ (64)
where ${\phi},\varphi,\gamma,\bar{\gamma},\Psi,\Phi,\lambda,\bar{\lambda}$ are defined in Appendix C. In particular, equations (C.12)–(C.15) for $\bar{\gamma},\bar{\lambda}$ are the solvable regulator equations because the output $y_{1}$ is a flat output of the considered plant (1)–(5). Besides, the functions ${\varsigma}(x,t)$ and $\bar{\varsigma}(x,t)$ satisfy
$\varsigma_{t}(x,t)=q_{2}{\varsigma}_{x}(x,t),~~\varsigma(0,t)=\frac{-1}{\vartheta(z_{1}(t),t)}f(\underline{z}_{n},t),$ (65)
$\bar{\varsigma}_{t}(x,t)=-q_{1}\bar{\varsigma}_{x}(x,t),~~\bar{\varsigma}(0,t)=p*\varsigma(0,t).$ (66)
The solution $\varsigma(x,t)$ is
$\varsigma(x,t)=\frac{-1}{\vartheta(z_{1}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}})}f\left(\underline{z}_{n}(t+\frac{x}{q_{2}}),t+\frac{x}{q_{2}}\right)$
@@
${\alpha_{t}}(x,t)=$ $-{q_{1}}{\alpha_{x}}(x,t),~~{\beta_{t}}(x,t)={q_{2}}{\beta_{x}}(x,t),$ (69)
$\alpha(0,t)=$ $p\beta(0,t),~~\beta(1,t)=0$ (70)
@@
$-(\acute{G}_{5}-\bar{\lambda}(1)){v}(t)+\varsigma(1,t)$ (71)
where $\varsigma(1,t)$ is given by (47)–(49), (54), (3.3), (B.5), ${\Psi}(1,y),{\Phi}(1,y)$, $\lambda(1)$ are determined by (C.16)–(C.21), and $\bar{\lambda}(1)$ is given by (C.12), (C.13), that is,
$\varsigma(1,t)=$ $\frac{-f(\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}{\vartheta(C_{1}\wp_{1}(Z(t),z[t],w[t],v(t),\frac{1}{q_{2}}),t+\frac{1}{q_{2}})}$
@@
Remark 2.
If actuator dynamics are included, the proposed safe output regulation scheme can be extended to an ODE–PDE–ODE configuration using the approach in [30]. It can also accommodate uncertain input delays in hyperbolic PDE–ODE cascades by integrating the method in [29].